### v2

@@ 5.2   Weighting scheme
The previous section provides the method to estimate each $CLATT_{e,e+l}$ in staggered DID-IV designs. In some applications, each $CLATT_{e,e+l}$ is of intrinsic interest if researchers are interested in the treatment effect heterogeneity across cohort $e$ and relative period $l$. However, in many cases, researchers may also want to estimate summary causal parameters, which capture the overall effect of a treatment on an outcome. Actually, in staggered DID-IV designs, each $CLATT_{e,e+l}$ depends not only on cohort $e$, relative period $l$, but also on the compliers $CM_{e,e+l}$. In this case, each $CLATT_{e,e+l}$ may be difficult to interpret when the number of cohorts and the relative period are fairly large, and when the members of compliers $CM_{e,e+l}$ in cohort $e$ vary and overlap over time.
@@
To propose the weighting functions for a variety of summary causal parameters in staggered DID-IV designs, we first define the average effect of the instrument on the treatment at a given relative period $l$ from the initial exposure to the instrument in cohort $e$. We call this the cohort specific average exposed effect on the treated in the first stage ($CAET_{e,e+l}^{1}$), formally defined below.
@@
In staggered DID designs, Callaway and Sant’Anna (2021) propose various aggregated measures along with different dimensions of treatment effect heterogeneity. We can employ their framework straightforwardly, but in staggered DID-IV designs, we should more carefully specify the weighting functions assigned to each summary measure.
@@
where the $CATT(\tilde{e},t)$ is the cohort specific average treatment effect on the treated at time $t$ in cohort $\tilde{e}$ (see, e.g., Callaway and Sant’Anna (2021), Sun and Abraham (2021)). Callaway and Sant’Anna (2021) construct the summary measure $\theta_{sel}(\tilde{e})$ by equally weighting each $CATT(\tilde{e},t)$ across all post-treatment periods because all the units in cohort $e$ adopt the treatment after they start exposed to that treatment.
@@
Notes: This table represents the specific expressions for the weights on each $CLATT(e,t)$ or each $CACRT(e,t)$ in a variety of summary causal parameters. Each target parameter corresponds with that of Callaway and Sant’Anna (2021). We superscript IV to associate each parameter with staggered DID-IV designs.
@@
where the weights are non-negative and sum to one by construction. This parameter summarizes each $CLATT_{\tilde{e},t}$ in cohort $\tilde{e}$ across all post-exposed periods and the weight assigned to each $CLATT_{\tilde{e},t}$ reflects the relative share of the compliers $CM_{\tilde{e},t}$ in period $t$ during the periods after the initial exposure to the instrument in cohort $\tilde{e}$. This weighting scheme would be reasonable in that it is designed to be larger at the period when the proportion of the compliers is relatively higher in cohort $\tilde{e}$.
By similar arguments, we can specify the weighting functions for various summary measures in staggered DID-IV designs, which correspond with those in Callaway and Sant’Anna (2021). In Table 7, we summarize the specific expressions for the weights assigned to each $CLATT_{\tilde{e},t}$ in each summary causal parameter. Here, we define each causal parameter by the notation analogous to that of Callaway and Sant’Anna (2021) and superscript IV to associate each parameter with staggered DID-IV designs. We note that our proposed weighting functions are the same if we have non-binary, ordered treatment where our target parameter is the $CACRT_{\tilde{e},t}$.
@@
We can construct the consistent estimator for the summary causal parameter $\theta^{IV}$ as follows:
@@
5.3   Pretrend test
In this section, we describe how one can conduct pretrend tests to check the validity of the parallel trends assumption in the outcome and the treatment.
Suppose that data are available for period $t=-1$. Then, one can assess the plausibility of Assumption 5 between period $t=-1$ and $t=0$ by testing the following null hypothesis
@@ 5.3   Pretrend test
We can generalize these tests to multiple pre-exposed periods settings by using the pretrends tests recently developed in the context of DID designs (Callaway and Sant’Anna (2021), Sun and Abraham (2021), Borusyak et al. (2021)); we can apply these tools to the first stage and the reduced form respectively to confirm the plausibility of Assumption 12 and 13.
@@
However, such tests are problematic if the effects of the instrument on the treatment and outcome are heterogeneous. Sun and Abraham (2021) shows that in staggered DID designs, the coefficient for pre-treatment periods on event study regressions can be different from zero even when the no anticipation assumption and the parallel trends assumption in the untreated outcome are plausible. Because we can view that $Y_{i,t}$ and $D_{i,t}$ are the outcomes and $Z_{i,t}$ is the binary absorbing treatment in equations (5)-(6), it is inappropriate to check whether the coefficients of pre-exposed periods are statistically different from zero or not as a way of testing Assumption 12 and Assumption 13.
@@ 6   Application
Oreopoulos (2006) estimates returns to schooling using a major education reform in the United Kingdom that increased the years of compulsory schooling from 14 to 15. Specifically, Oreopoulos (2006) exploit variation resulting from the different timing of implementation of school reforms between Britain (England, Scotland, and Wales) and Northern Ireland as an instrument for education attainment: the school-leaving age increased in Britain in 1947, while the reform was not implemented until 1957 in Northern Ireland. The data are a sample of individuals in Britain and Northern Ireland, who were aged $14$ between $1936$ and $1965$, constructed from combining the series of U.K. General Household Surveys between $1984$ and $2006$; see Oreopoulos (2006), Oreopoulos (2008) for details.
In the first part of their analysis, Oreopoulos (2006) adopts regression discontinuity designs (RDD) and analyzes the data sets in Britain and Northern Ireland separately. Due to the imprecision of their standard errors [^7] Kolesár and Rothe (2018) consider inference in regression discontinuity designs with a discrete running variable and apply their theory to Oreopoulos (2006)., Oreopoulos (2006) then moves to ”a difference-in-differences and instrumental-variables analysis by combining the two sets of U.K. data” to estimate some local average treatment effects.
Specifically, Oreopoulos (2006) (more precisely, Oreopoulos (2008)) runs the two-way fixed effects instrumental variable regression, using the reform as an excluded instrument for education attainment:
@@
Here, a cohort (a year when aged $14$) plays a role of time as it determines the exposure to the policy change. The dependent variables $Y_{i,t}$ and $D_{i,t}$ are log annual earnings and education attainment for unit $i$ and cohort $t$ respectively. Both the first stage and the reduced form regressions include a birth cohort fixed effect and a North Ireland fixed effect. The binary instrument $Z_{i,t}\in\{0,1\}$ takes one if unit $i$ in cohort $t$ is exposed to the policy shock.
The staggered introduction of the school reform can be viewed as a natural experiment, but is not randomized in reality. Oreopoulos (2006) notes that the reform was implemented with political support, taking into account the cost and benefit of the school reform. This indicates that Oreopoulos (2006) implicitly relies on a staggered DID-IV identification strategy instead of the random variation of the policy change. Actually, Oreopoulos (2006) presents the corresponding plots of British and Northern Irish average education attainment and average log earnings by cohort to illustrate the evolution of these variables before and after the policy shock.
In the author’s setting, however, the effect of the school reform on education attainment and log annual earnings can differ between Britain and North Ireland, and potentially vary across cohorts. The TWFEIV estimate, therefore, can potentially fail to aggregate the causal effects and yield a misleading conclusion.
@@ 6.1   Assessing the identifying assumptions in staggered DID-IV designs
We first discuss the validity of the staggered DID-IV identification strategy in Oreopoulos (2006). In the author’s setting, our target parameter is the cohort specific average causal response on the treated (CACRT) as education attainment is a non-binary, ordered treatment. We can identify each CACRT and its summary measures if the underlying DID-IV identifying assumptions seem plausible.
@@
It would be automatically satisfied in the author’s setting: the policy reform (instrument) increased the minimum schooling-leaving age from $14$ to $15$, which ensures that there are no defiers during the periods after the policy shock.
@@
It would be plausible that there is no anticipation before the policy reform, if the treatment adoption behavior is the same as the one in the absence of the policy change before its implementation in England. This assumption may be violated if units have private knowledge about the probability of extended education and manipulate their education attainment before the policy shock.
Next, we assess the validity of the parallel trends assumptions in the treatment and the outcome. To do so, we use the interacted two-way fixed effects regressions proposed by Sun and Abraham (2021) in the first stage and the reduced form, respectively. The results are shown in Figure 2.
Figure 2: The effect of the instrument in the first stage and reduced form in the setting of Oreopoulos (2006). Notes: The results for the effect of the school reform on education attainment (Panel (a)) and on log annual earnings (Panel (b)) under the staggered DID-IV identification strategy. The unexposed group $U$ is a last-exposed cohort, North Ireland and the reference period is $t=1946$. The blue lines represent the estimates with pointwise $95\%$ confidence intervals for pre-exposed periods in both panels. These should be equal to zero under the null hypothesis that the parallel trends assumptions in the treatment and the outcome hold. The red lines represent the estimates with pointwise $95\%$ confidence intervals for post-exposed periods in both panels.
@@
It requires that the expectation of log annual earnings would have followed the same evolution between England and North Ireland across cohorts if the reform had not occurred. Panel (b) in Figure 2 plots the results of the interacted two-way fixed effects regression in the reduced form along with a $95\%$ pointwise confidence interval. The pre-exposed estimates seems consistent with Assumption 13: though there exists a downward pretrends, the estimates before the exposure to the policy change are not significantly different from zero.
Figure 2 also sheds light on the dynamic effects of the policy reform on education attainment and log annual earnings in post exposed periods. In Panel (a), the estimated increase in years of schooling after the reform ranges from $0.56$ to $0.87$, and all the estimates are statistically significant. In Panel (b), the estimated increase in log annual earnings after the reform varies from $8\%$ to $25\%$, and $5$ out of $10$ estimates are statistically significant.
We note that the post-exposed estimates in Panel (b) do not capture each CACRT in post-exposed periods: each estimate in the reduced form is not scaled by the corresponding estimate in the first stage. In the next section, we estimate each CACRT in post-exposed periods by employing our proposed estimation method.
@@ 6.2   Illustrating our estimation method
We start by estimating two-way fixed effects instrumental variable regression in the author’s setting. In our estimation, we slightly modify the author’s specification to clearly illustrate the pitfalls of TWFEIV regression. Specifically, Oreopoulos (2006) (more precisely Oreopoulos (2008)) includes some covariates (survey year, sex, and a quartic in age) and runs the weighted regression in the main specification (see Oreopoulos (2006), Oreopoulos (2008) for details), whereas we exclude such covariates and do not apply their weights to our regression.
The result is shown in Table 8. The TWFEIV estimate is $-0.009$ and it is not significantly different from 00. This implies that on the whole, the returns to schooling in the U.K. may be nearly zero. However, this may be a misleading conclusion: we cannot interpret that the TWFEIV estimand captures the properly weighted average of each CACRT if the effects of the policy reform on education attainment and log annual earnings are not stable across cohorts (Miyaji (2024)).
The TWFEIV regression performs the ”bad comparisons” (c.f. Goodman-Bacon (2021)) in the author’s setting: when we run the TWFEIV regression, we treat the set of already exposed units in North Ireland as controls between $1957$ and $1965$ because both regions are already exposed to the policy shock during these periods. In Appendix D, we quantify the bias terms of the TWFEIV estimand by using Lemma 7 in Miyaji (2024), and show that the TWFEIV estimand is negatively biased in Oreopoulos (2006). Specifically, we show that all the bias terms are positive, whereas all the assigned weights are negative, which yields the downward bias for the TWFEIV estimand.
Figure 3: The cohort specific average causal response on the treated in each relative period in the setting of Oreopoulos (2006). Notes: The results for returns to schooling under the staggered DID-IV identification strategy. The unexposed group $U$ is North Ireland and the reference period is period $t=1946$. The red lines represent the stacked two stage least squares estimates with pointwise $95\%$ confidence intervals for post-exposed periods.
We now employ our proposed method to estimate each CACRT. We first create data sets by cohort $t$ ($t=1947,\dots,1956$). Each data set only contains units of cohort $t$ and cohort $t=1946$ in England and North Ireland. We define North Ireland as an unexposed group $U$ and discard the units who were aged at $14$ between $1957$ and $1965$ in order to avoid the bad comparisons. We then run TSLS regression in each data set. The standard error is calculated by using the influence function shown in equation (C.2) in Appendix C.
@@
Finally, we estimate the summary causal measure constructed from each $CACRT_{e,t}$. Specifically, we estimate the weighted average of each $CACRT_{e,t}$ during post-exposed periods in England ($e=1947$):
@@
Here, each weight assigned to each $CACRT_{e,t}$ represents the relative amount of the effect of the policy reform on education attainment during post-exposed periods in England.
The result is shown in Table 8. The estimate is $0.24$ and it is significantly different from zero. The estimated returns to schooling are substantial: this is likely because each $CACRT_{e,t}$ captures the returns to schooling among the compliers: such units may belong to relatively low-skilled labor or low-income family, and have potentially much gain from the school reform.
The estimates obtained from our proposed method and weighting scheme are significantly different from the TWFEIV estimate: our STS estimates and its summary measure are all positive, whereas the TWFEIV estimate is strictly negative. Overall, our results indicate that the economic returns of education are substantial in the U.K. and the estimation method matters in staggered DID-IV designs in practice.

### v4

@@ 5.2   Weighting scheme
The previous subsection provides the method to estimate each $CLATT_{e,e+l}$ in staggered DID-IV designs. In some applications, each $CLATT_{e,e+l}$ is of intrinsic interest if researchers are interested in the treatment effect heterogeneity across cohort $e$ and relative period $l$. However, researchers may also want to estimate summary causal parameters that capture the overall effect of a treatment on an outcome. In staggered DID-IV designs, each $CLATT_{e,e+l}$ depends not only on cohort $e$, relative period $l$, but also on the compliers $CM_{e,e+l}$. In this case, each $CLATT_{e,e+l}$ may be difficult to interpret when the number of cohorts and the relative period are fairly large, and when the members of compliers $CM_{e,e+l}$ in cohort $e$ vary and overlap over time.
@@
To propose the weighting functions for a variety of summary causal parameters in staggered DID-IV designs, we define the average effect of the instrument on the treatment at a given relative period $l$ from the initial exposure to the instrument in cohort $e$, called the cohort specific average exposed effect on the treated in the first stage ($CAET_{e,e+l}^{1}$).
@@
In staggered DID designs, Callaway and Sant’Anna (2021) propose various aggregated measures along with different dimensions of treatment effect heterogeneity. We can employ their framework directly, but in staggered DID-IV designs, we should more carefully specify the weighting functions assigned to each summary measure.
@@
where $CATT(\tilde{e},t)$ is the cohort specific average treatment effect on the treated at time $t$ in cohort $\tilde{e}$ (see, e.g., Callaway and Sant’Anna (2021), Sun and Abraham (2021)). Callaway and Sant’Anna (2021) construct the summary measure $\theta_{sel}(\tilde{e})$ by equally weighting each $CATT(\tilde{e},t)$ across all post-treatment periods because all the units in cohort $e$ adopt the treatment after they start being exposed to it.
@@
Notes: This table represents the specific expressions for the weights on each $CLATT(e,t)$ or each $CACRT(e,t)$ in a variety of summary causal parameters. Each target parameter corresponds with that of Callaway and Sant’Anna (2021). IV is added to associate each parameter with staggered DID-IV designs.
@@
where the weights are non-negative and sum to one by construction. This parameter summarizes each $CLATT_{\tilde{e},t}$ in cohort $\tilde{e}$ across all post-exposed periods and the weight assigned to each $CLATT_{\tilde{e},t}$ reflects the relative share of the compliers $CM_{\tilde{e},t}$ in period $t$ during the periods after the initial exposure to the instrument in cohort $\tilde{e}$. This weighting scheme would be reasonable in that it is designed to be larger in the period when the proportion of the compliers is relatively higher in cohort $\tilde{e}$.
By similar arguments, we can specify the weighting functions for various summary measures in staggered DID-IV designs, which correspond with those in Callaway and Sant’Anna (2021). In Table 7, we summarize the specific expressions for the weights assigned to each $CLATT_{\tilde{e},t}$ in each summary causal parameter. Here, we define each causal parameter by the notation analogous to that of Callaway and Sant’Anna (2021) and superscript IV to associate each parameter with staggered DID-IV designs. Note that our proposed weighting functions are the same under non-binary, ordered treatment settings, in which our target parameter is the $CACRT_{\tilde{e},t}$.
@@
We can construct the consistent estimator for the summary causal parameter $\theta^{IV}$ as follows.
@@
5.3   Pre-trend test
We now describe how one can conduct pre-trend tests to check the validity of the parallel trends assumption in the outcome and the treatment.
Suppose that data are available for period $t=-1$; one can assess the plausibility of Assumption 5 between period $t=-1$ and $t=0$ by testing the following null hypothesis
@@ 5.3   Pre-trend test
These tests can be generalized to multiple pre-exposed periods settings by using the pre-trend tests recently developed in the context of DID designs (Callaway and Sant’Anna (2021), Sun and Abraham (2021), Borusyak et al. (2021)); we can apply these tools to the first stage and the reduced form respectively to confirm the plausibility of Assumption 12 and 13.
@@
However, such tests are problematic if the effects of the instrument on the treatment and outcome are heterogeneous. Sun and Abraham (2021) shows that in staggered DID designs, the coefficient for pre-treatment periods on event study regressions can be different from zero even when the no anticipation assumption and parallel trends assumption in the untreated outcome are plausible. Because we can view that $Y_{i,t}$ and $D_{i,t}$ are the outcomes and $Z_{i,t}$ is the binary absorbing treatment in equations (5)-(6), it is inappropriate to check whether the coefficients of pre-exposed periods are statistically different from zero or not as a way of testing Assumption 12 and Assumption 13.
@@ 6   Application
Oreopoulos (2006) estimates returns to schooling using a major education reform in the UK that increased the years of compulsory schooling from 14 to 15. Specifically, Oreopoulos (2006) exploit variation resulting from the different timing of implementation of school reforms between Britain (England, Scotland, and Wales) and Northern Ireland as an instrument for education attainment: the school-leaving age increased in Britain in 1947, but was not implemented until 1957 in Northern Ireland. The data are a sample of individuals in Britain and Northern Ireland, who were aged $14$ between $1936$ and $1965$, constructed from combining the series of U.K. General Household Surveys between $1984$ and $2006$; see Oreopoulos (2006), Oreopoulos (2008) for details.
In the first part of his analysis, Oreopoulos (2006) adopts regression discontinuity designs (RDD) and analyzes the data sets in Britain and Northern Ireland separately. Due to the imprecision of their standard errors [^8] Kolesár and Rothe (2018) consider inference in regression discontinuity designs with a discrete running variable and apply their theory to Oreopoulos (2006)., Oreopoulos (2006) then moves to ”a difference-in-differences and instrumental-variables analysis by combining the two sets of U.K. data”.
Specifically, Oreopoulos (2006) (more precisely, Oreopoulos (2008)) runs the following two-way fixed effects instrumental variable regression with the education reform as an excluded instrument for education attainment.
@@
Here, a cohort (a year when aged $14$) plays a role of time as it determines the exposure to the policy change. The dependent variables $Y_{i,t}$ and $D_{i,t}$ are log annual earnings and education attainment for unit $i$ and cohort $t$, respectively. Both the first stage and the reduced form regressions include a birth cohort fixed effect and a North Ireland fixed effect. The binary instrument $Z_{i,t}\in\{0,1\}$ takes one if unit $i$ in cohort $t$ is exposed to the policy change.
The staggered introduction of the school reform can be viewed as a natural experiment, but is not randomized across regions in reality; Oreopoulos (2006) notes that the reform was implemented with political support, taking into account costs and benefit. This indicates that Oreopoulos (2006) implicitly relies on a staggered DID-IV identification strategy instead of the IV design. Indeed, Oreopoulos (2006) presents the corresponding plots of British and Northern Irish average education attainment and average log earnings by cohort to illustrate the evolution of these variables before and after the policy shock.
In the author’s setting, however, the effect of the school reform on education attainment and log annual earnings can differ between Britain and North Ireland, potentially varying across cohorts. That the author was conscious of the heterogeneous treatment effects is clear from the title of ”Estimating Average and Local Average Treatment Effects of Education when Compulsory Schooling Laws Really Matter”. The TWFEIV estimator, therefore, can potentially fail to aggregate the treatment effects, yielding a misleading conclusion under staggered DID-IV designs.
@@ 6.1   Assessing the identifying assumptions in staggered DID-IV designs
We discuss the validity of the staggered DID-IV identification strategy in Oreopoulos (2006). In the author’s setting, our target parameter is the cohort specific average causal response on the treated (CACRT), as education attainment is a non-binary, ordered treatment. We can identify each CACRT and its summary measure if the underlying DID-IV identifying assumptions seem plausible.
@@
It would be automatically satisfied in the author’s setting: the policy change (instrument) increased the minimum schooling-leaving age from $14$ to $15$, ensuring that there are no defiers during the periods after the policy shock.
@@
It would be plausible that there is no anticipation, if the treatment adoption behavior is the same as the one in the absence of the policy change before its implementation in England. This assumption may be violated if units have private knowledge about the probability of extended education and manipulate their education attainment before the policy shock.
Next, we assess the validity of the parallel trends assumptions in the treatment and the outcome, using the interacted two-way fixed effects regressions proposed by Sun and Abraham (2021) in the first stage and the reduced form, respectively. The results are shown in Figure 2.
Figure 2: The effect of the instrument in the first stage and reduced form in the setting of Oreopoulos (2006). Notes: This figure presents the results of the effect of the school reform on education attainment (Panel (a)) and on log annual earnings (Panel (b)) under the staggered DID-IV identification strategy. The unexposed group $U$ is a last-exposed cohort, North Ireland and the reference period is $t=1946$. The blue lines represent the estimates with pointwise $95\%$ confidence intervals for pre-exposed periods in both panels. These should be equal to zero under the null hypothesis that the parallel trends assumptions in the treatment and the outcome hold. The red lines represent the estimates with pointwise $95\%$ confidence intervals for post-exposed periods in both panels.
@@
It requires that the expectation of log annual earnings would have followed the same evolution between England and North Ireland across cohorts in the absence of the school reform. Panel (b) in Figure 2 plots the results of the interacted two-way fixed effects regression in the reduced form along with a $95\%$ pointwise confidence interval. The pre-exposed estimates seems consistent with Assumption 13: though an upward pre-trend exists, all the estimates before the initial exposure to the policy change are not significantly different from zero.
Figure 2 also sheds light on the dynamic effects of the policy reform on education attainment and log annual earnings in post-exposed periods. In Panel (a), the estimated increase in years of schooling after the reform ranges from $0.56$ to $0.87$, and all estimates are statistically significant. In Panel (b), the estimated increase in log annual earnings after the reform varies from $8\%$ to $25\%$, and $5$ out of $10$ estimates are statistically significant.
Note that the post-exposed estimates in Panel (b) do not capture each CACRT in post-exposed periods: each estimate in the reduced form is not scaled by the corresponding estimate in the first stage. In the next section, we estimate each CACRT in post-exposed periods by employing our proposed estimation method.
@@ 6.2   Illustrating our estimation method
We start by estimating two-way fixed effects instrumental variable regression in the author’s setting. To clearly illustrate the pitfalls of TWFEIV regression, in our estimation, we slightly modify the author’s specification. Oreopoulos (2006) (more precisely Oreopoulos (2008)) includes some covariates (survey year, sex, and a quartic in age) and runs the weighted regression in the main specification (see Oreopoulos (2006), Oreopoulos (2008) for details), whereas we exclude such covariates and do not apply their weights to our regression.
The result is shown in Table 8. The TWFEIV estimate is $-0.009$, not significantly different from 00, which indicates that, on the whole, the returns to schooling in the UK is nearly zero. However, this may be a misleading conclusion. We cannot interpret that the TWFEIV estimand captures the properly weighted average of all possible CACRTs if the effect of the school reform on education attainment or log annual earnings is not stable across cohorts (Miyaji (2024)).
The TWFEIV regression performs the ”bad comparisons” (c.f. Goodman-Bacon (2021)) in the author’s setting. When we run the TWFEIV regression, we treat the set of already exposed units in North Ireland as controls between $1957$ and $1965$, as both regions are already exposed to the policy shock during these periods. In Appendix D, we quantify the bias terms of the TWFEIV estimand by using Lemma 7 in Miyaji (2024), and show that the TWFEIV estimand is negatively biased in Oreopoulos (2006). Specifically, we show that all the bias terms are positive, whereas all the assigned weights are negative, which yields the downward bias for the TWFEIV estimand.
Figure 3: The cohort specific average causal response on the treated in each relative period in the setting of Oreopoulos (2006). Notes: This figure shows the results for returns to schooling under the staggered DID-IV identification strategy. The unexposed group $U$ is North Ireland and the reference period is period $t=1946$. The red lines represent the stacked two stage least squares estimates with pointwise $95\%$ confidence intervals for post-exposed periods.
We now employ the proposed method to estimate each CACRT. First, we create data sets by cohort $t$ ($t=1947,\dots,1956$). Each data set only contains units of cohort $t$ and cohort $t=1946$ in England and North Ireland. We define North Ireland as an unexposed group $U$ and discard the units aged at $14$ between $1957$ and $1965$ to avoid the bad comparisons. We then run TSLS regression in each data set. The standard error is calculated by using the influence function shown in equation (C.2) in Appendix C.
@@
Finally, we estimate the summary causal measure by aggregating each $CACRT_{e,t}$. Specifically, we estimate the weighted average of each $CACRT_{e,t}$ during post-exposed periods in England ($e=1947$).
@@
Here, each weight assigned to $CACRT_{e,t}$ represents the relative amount of the effect of the policy reform on education attainment during post-exposed periods in England.
The result is shown in Table 8. The estimate is $0.24$ and it is significantly different from zero. The estimated returns to schooling are substantial, perhaps because each $CACRT_{e,t}$ captures the returns to schooling among the compliers: such units may belong to relatively low-skilled labor or low-income family that potentially have much to gain from school reform.
The estimates obtained from our proposed method and weighting scheme are significantly different from the TWFEIV estimate: our STS estimates and its summary measure are all positive, whereas the TWFEIV estimate is strictly negative. Overall, our results indicate that the economic returns of education are substantial in the UK, and the estimation method matters in staggered DID-IV designs in practice.